OpenStudy (anonymous):

determine whether the given poin is a solution of 2y=-12x+6 (-1/2,6)

OpenStudy (anonymous):

So to figure this out just plug in for x and y 2y=-12x+6 2(6)=-12(-1/2)+6 12=6+6 12=12 Yes it is a solution